Let S = {1, 2, 3, 4}. The total number of unordered pairs of disjoint subsets of S is equal to?

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
GivenS = {1, 2, 3, 4}.
So it consists of 4 elements.
Asked to find the total number of unordered pairs of disjoint subsets of S.
So number of unordered pairs=(34+1)/2=82/2=41